OpenStudy (anonymous):

double check my answer? The product of (a − b)(a − b) is a^2 − b^2. The choices are always, sometimes, and never. I chose always is that correct?

OpenStudy (anonymous):

thanks(:

OpenStudy (anonymous):

erm...The product \[(a − b)(a − b) = a^2 -2ab + b^2\]

OpenStudy (anonymous):

so the answer is not always?

OpenStudy (anonymous):

correct, it would only be that \[(a−b)(a−b)=a^2+b^2\] if either a or b is zero. In that case, you would get either \[a^2\] or \[b^2\] only. So sometimes true i.e. only when a=0 or b=0 or both (trivial). If a and b are each not zero, then it is not true.

OpenStudy (anonymous):

i guess it could be true of b was zero

OpenStudy (anonymous):

it would not be true if \[a=0\]

OpenStudy (anonymous):

@satellite73 it true for both man , when it is a=0 0r b=0

OpenStudy (mathmate):

Sometimes, (a-b)^2=a^2-b^2 iff b=0 if a=0, b not equal to zero, then (a-b)^2=+b^2 a^2-b^2=-b^2